OpenStudy (anonymous):

Simplify -11sqrt 112

OpenStudy (anonymous):

Break the argument of the square root into a prime factorization, then pull out all the perfect squares from inside.\[-11\sqrt {112}=-11\sqrt{2^4*7}=-44\sqrt{7}\]

OpenStudy (campbell_st):

multiply by \[\sqrt{112}/\sqrt{112}\] gives \[-11 \times \sqrt{112}/112\\[= -11 \times4\sqrt{7}/112\]] \[= -11\sqrt{7}/28\]

OpenStudy (anonymous):

Campbell, one of us didn't understand the question correctly.

OpenStudy (anonymous):

Can you please help me I am taking a test and the problem above was not a given answer!

OpenStudy (anonymous):

Did you look at my answer?

OpenStudy (anonymous):

See it now thank you